OpenStudy (anonymous):

A nuclear power plant has a power output of 1000MW and operates with an efficiency of 33%. If excess hear is carried away from the plant by a river with a flow rate of 1.0x10^6 kg/s, what's the rise in temperature of the flowing water?

OpenStudy (anonymous):

Efficiency of 33% implies 67% is rejected as heat. To get 1000MW output 3030MW will need to be input. Assuming waste energy is heat and is all transported through the cooling water, 2030 MJ of heat energy will be injected into 1E6 kg of water every second. Using a specific heat capacity of 4181.3 J/kg/K (@25 C) gives delta T = E / (m.C) = 2030E6 / (1E6 . 4181.3) or 0.486 K